2√ [48] + √[27] - 5√[12]
I first looked at 2√[48]which tells me it's the square root of 48 so I then broke 48 down into factors.
48 = 2 * 2 * 2 * 2 * 3
I took out 2 * 2 = 4 and left 3 under the radical sign 4√[3]
I the did the same for √[27] following the same steps and I got 3√[3]. where my problem comes in is when I get to 5√[12] how do I break this down. I also wanted to know am I doing this correctly. PLease point me in the right direction. I am having troubles seeing how to break these down.

Answer by checkley71(8403)

SQRT48=SQRT(2*2)(2*2)3 OR 2*2SQRT3 OR 4SQRT3
SQRT27=SQRT(3*3)3 OR 3SQRT3
SQRT12=SQRT(2*2)3 OR 2SQRT3